` HUBBLE SPACE TELESCOPE DAILY REPORT #2265 PERIOD COVERED: 0000Z (UTC) 12/07/98 - 0000Z (UTC) 12/08/98 Daily Status Report as of 342/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Thirteen Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.3 Completed STIS/CCD/MA2 7400 (Evidence from Boron for Massive Star Evolution) ] The Space Telescope Imaging Telescope (CCD/MA2) was used to observe the B III and B II resonance lines at 2065.8 and 1362.5 Angstroms, respectively, in the main-sequence and slightly evolved blue star PG0832+676 using the HST STIS MAMA detectors and echelle gratings. There was a Guide Star Acquisition (1,3,3) which resulted in a Fine Lock backup on FGS#3 at 342/1015Z. HSTAR #6803 was written. The proposal completed with no other anomalies reported. 1.4 Completed Two Sets of WF/PC-2 6362 (Supernova Host Galaxy) The WF/PC-2 was used to make observations of the supernova host galaxies GAL-33638 and GAl-31228. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 1.6 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxy 3C428.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.7 Completed Three Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.8 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy VCC1822 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.9 Completed WF/PC-2 7409 (Resolving sdB Binary Systems) The WF/PC-2 was used to make observations to try to resolve the sdB class binary system STAR-22162. This is a continuation of the approved Cycle 6 Proposal No. 6351, which due to unlucky circumstances did not succeed in obtaining any data until 8/30/96. The observations were completed as scheduled, and no anomalies were reported. 1.10 Completed Three Sets of WF/PC-2 6363 (The Environments and Host Galaxies of BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of BL-Lac objects BL-Lacerta, MS0922.9+7 AND PKS0235+16 in order to study their environment and their host galaxies. The goal is to investigate systematic trends in host galaxy type, unusual morphological features, nearby companion galaxies, etc. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 7514 (HST Observations of Orion -- Probing the Origin of Abundance Anomalies in H II Regions) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the H-II region (i.e., NGC-1976) which is within the Orion Nebula (a.k.a., M-42) in order to probe the origin of the abundance anomalies in the nebula. The purpose the of observations is to obtain a set of emission-line diagnostics of the Orion Nebula in its brightest and best-resolved H II region. A calibration Wavelength image and CCD flat were also taken. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed Three Sets of WF/PC-2 6666 (The Mass Function and Binary Star Fraction of the Trapezium Cluster) The WF/PC-2 was used to take deep images of the Trapezium cluster in Orion to determine the low-mass end of the initial mass function, to search for brown dwarfs, and to explore the close binary population in the cluster. The observations completed as planned with no problems. 1.13 Completed STIS/CCD 8081 (STIS CCD Hot Pixel Annealing -- II) The Space Telescope Imaging Spectrograph (CCD) was used to make a dark and a bias calibration observation. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 6633 (Cusps or Cores in Barred Bulges?) The WF/PC-2 was used to make observations of the SB class (Barred) galaxy NGC-2633 in order to study its nuclear bulge region and cusps. This is a highly tri-axial, tumbling, low-mass stellar systems. The observations were completed as planned, and no problems were reported. 1.15 Completed WF/PC-2 6359 (Core Properties of the Bulges of Spiral Galaxies) The WF/PC-2 was used to make snapshot survey observations of spiral galaxy NGC-7188 in order to study their central bulges. The aim is to investigate the core properties, correlations with global galactic properties, and the frequency of occurrence of multiple and/or miscentered nuclei. The observations were completed as scheduled, and no anomalies were noted. 1.16 Completed WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.17 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.18 Completed NIC/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Cameras Three) was used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 1 Successful : 1 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: There were a total of six IP Data hits. TTR 20433 was written. The SSR EDAC Error counter was cleared two times. (ROP SR-1). There was a Guide Star Acquisition (1,3,3) which resulted in a Fine Lock backup on FGS #3 at 342/1015Z. See section 1.3 for affected proposal (7400). HSTAR #6803 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH